# Efficient 3D Molecular Generation with Flow Matching and Scale Optimal Transport

### v2

@@ 6 Related Work
In addition to the unconditional molecular generators we outlined above, a number of works have attempted to directly generate ligands within protein pockets [35, 15, 42]. However, these models also suffer from the issues we outlined previously, including long-sampling times (100s - 1000s of seconds for 100 molecules [42]) and generating invalid chemical structures or molecules with very high strain energies [17]. GraphBP [29], an autoregressive model for protein-conditioned generation, is able to generate ligands faster, but suffers from worse docking scores than more recent diffusion models.
@@
Outside of small molecule design flow-matching has recently gained traction with generative models for biomolecules. FoldFlow [8] and FrameFlow [60] are both recently introduced flow-matching models for protein structure generation. Multiflow [10] attempts to jointly generate protein sequence and structure and introduces the discrete flow models (DFM) framework for flow-matching generation of discrete data. Stark et al. [49] also introduce a framework for flow-matching on discrete data, DirichletFM, and apply this to DNA sequence design. Finally, Verma et al. [54] use a conjoined system of ODEs to train a model to jointly generate antibody sequences and structures.
@@ 7 Conclusion
In this work we have presented Semla, a novel equivariant message passing architecture which exhibits much better efficiency and scalability than existing approaches for unconditional molecular generation. Additionally, we introduced a novel type of optimal transport (OT), scale OT, which extends equivariant OT but noticeably reduces the transport cost. We then trained a Semla model for 3D molecular generation using flow matching with scale OT. Our model, SemlaFlow, achieves state-of-the-art results on molecular generation benchmarks and, crucially, is able to generate high quality molecules with only 20 sampling steps. We also highlighted issues with current molecular evaluation metrics and proposed the use of energy and strain for evaluating the quality of generated molecular conformations.
@@
While we believe our model has made significant progress in solving key challenges for 3D molecular generators, many challenges remain. Firstly, the energies of the molecules generated by SemlaFlow are still significantly higher than that of the dataset; generating molecular coordinates with very high fidelity remains a problem for these models. Including further inductive biases or fine-tuning against an energy model could be an avenue to improve this in future work [34, 43, 57]. Additionally, while SemlaFlow has shown significant efficiency improvements over existing methods, it still uses a full-connected message passing component, limiting it’s scalability to larger molecular systems. We leave the further enhancement of the scalability of Semla to future work. We believe our model makes crucial step towards the practical application of 3D molecular generators, although we leave the integration of SemlaFlow into drug discovery workflows, either through RL-based fine-tuning or protein pocket conditioned generation, to future work. In its current form our model has the potential to be used for malicious activities, however, we believe the potential positive contributions of this model, helping to design new therapeutics, for example, strongly outweigh the negatives.
Acknowledgments and Disclosure of Funding
@@
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Preliminary experiments were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS, project: 2024/22-33), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
The authors thank T. Le (Pfizer) for sharing code and weights of [26] ahead of publication.
Appendix A Scale Optimal Transport
Figure 3 shows the transport cost for different sizes of molecules. We measure the transport cost as the mean-squared error between a batch of molecular coordinates and a batch of coordinate noise sampled using either equivariant OT or scale OT. For each batch of molecules we sample 10 different batches of noise and calculate the mean transport cost. We use a batch of 100 as a representative example. Although the distribution is much wider, we only show transport cost values for molecules with between 20 and 70 atoms since almost all of the probability mass lies within this region. A histogram of the frequency of different sizes of molecules is also shown in Figure 3. We can see that scale OT provides a significant reduction in the transport cost overall. The reduction is very pronounced for smaller molecules but becomes less so as the molecular size increases. Even for molecules with between 40 and 50 atoms – the densest part of the distribution – scale OT still provides a noticeable reduction in the transport cost.
Figure 3: Left: average transport cost (mean squared error) for molecules of different sizes with both equivariant and scale optimal transport. Shaded areas show 2 standard deviations either side of the mean. Right: histogram of the number of atoms per molecule for the GEOM Drugs dataset.
Appendix B Ablation Experiments
In Table 4 we present the results of ablation experiments showing the importance of various model components. The results in the table work in the following way – Baseline refers to a SemlaFlow model with one coordinate set, using equivariant optimal transport and no self conditioning. The models below then build on each other, starting at the top and working down the table, adding the component given in the model column. For example, the model in the row marked by + Scale OT uses 64 coordinate sets and scale optimal transport. The final model is the same as presented in the main text but trained for only 100 epochs, as with the other models in Table 4.
The ablation results show that all three additions are important for model performance. Unlike using additional coordinate sets and self conditioning, scale optimal transport comes with no increase in training time. Scale OT leads to a significant reduction in the energies and strains of the generated molecules, although does come at the cost of slightly decreased validity. In practice we believe this is still a very beneficial trade-off – validity becomes much less important when it’s close to 100% and the lower energies and strains show that the model is producing higher quality conformations.
Table 4: SemlaFlow ablation results on the Geom Drugs dataset. Each model was trained for 100 epochs. During evaluation an average over 5 evaluation runs was taken where each run consisted of sampling 5000 molecules with sizes sampled randomly from the test set.
Model Mol Stab Validity Energy Strain Uniqueness Novelty
Baseline 88.8±${}_{\pm\text{0.54}}$ 82.7±${}_{\pm\text{0.43}}$ 692.0±${}_{\pm\text{2}6.5}$ 662.3±${}_{\pm\text{26.4}}$ 100.0±${}_{\pm\text{0.0}}$ 99.7±${}_{\pm\text{0.03}}$
+ Coord Sets 96.2±${}_{\pm\text{0.11}}$ 95.0±${}_{\pm\text{0.12}}$ 318.7±${}_{\pm\text{4.08}}$ 269.3±${}_{\pm\text{3.48}}$ 99.9±${}_{\pm\text{0.01}}$ 99.8±${}_{\pm\text{0.06}}$
+ Scale OT 95.9±${}_{\pm\text{0.23}}$ 91.9±${}_{\pm\text{0.14}}$ 199.5±${}_{\pm\text{2.12}}$ 161.4±${}_{\pm\text{2.12}}$ 100.0±${}_{\pm\text{0.0}}$ 99.7±${}_{\pm\text{0.08}}$
+ Self Cond 97.4±${}_{\pm\text{0.18}}$ 94.6±${}_{\pm\text{0.24}}$ 135.9±${}_{\pm\text{2.0}}$ 98.4±${}_{\pm\text{2.30}}$ 100.0±${}_{\pm\text{0.0}}$ 99.7±${}_{\pm\text{0.06}}$
Appendix C Training Details
All SemlaFlow models presented in this paper are constructed from 12 Semla layers, with $D=512$, $M=128$ and $S=64$, where $D$ is the model dimension, $M$ is the message dimension, and $S$ is the number of coordinate sets. This corresponds to approximately 40M learnable parameters, which is more than 3 times as many than the current state-of-the-art, EQGAT-diff [26]. Despite this additional expressivity, our model processes batches more than twice as quickly due to the efficiency improvements from Semla.
@@ Appendix C Training Details
All models were trained with the AMSGrad [38] variant of the Adam optimiser [22] with a learning rate (LR) of 0.0003. We also apply linear LR warm-up, using 2000 warm-up steps for QM9 and 10000 warm-up steps for GEOM Drugs. During training we clip the norms of the gradients at 1.0 for all models. Loss weightings $(\lambda_{x},\lambda_{a},\lambda_{b},\lambda_{c})=(1.0,0.2,0.5,1.0)$ were used to train QM9 models. The same weightings were used for GEOM Drugs, except we set $\lambda_{b}=1.0$.
@@
In order to make the training as efficient as possible we place molecules in the dataset into buckets based on their size, and then form minibatches for training and evaluation within the buckets. This ensures that all batches contain similarly sized molecules so that the amount of padding within each batch is minimised. With this setup we can also apply a cost function to select the batch size for each bucket separately; since the amount of memory required to process a molecule increases quadratically with the number of atoms, this helps to balance the GPU memory consumption for each batch. In practice, though, we simply apply a linear cost function and use a batch size of 4096 atoms per batch for all our models. While we have found our bucketing scheme leads to a significant increase in training speed, it may also introduce additional bias into the training since molecules within each batch are no longer selected completely at random. Although we have not attempted to quantify this bias our results seem to show that bucketing is not significantly detrimental to performance.
Appendix D Issues with Current Evaluation Metrics
@@
  • Validity is given by the proportion of molecules which can be successfully sanitized using RDKit.
  • Uniqueness measures the proportion of generated molecules which are unique in their 1D (i.e. SMILES string) form.
  • Novelty measures the proportion of molecules which are not in the training set.
While conducting evaluations on SemlaFlow and existing models we have noticed a number of problems with these metrics, which we outline as follows:
  1. 1.
    Different models have used different lookup tables to compute atom and molecule stability, potentially leading to different evaluation results on the same molecules. We have found that this is mostly due to whether models predict formal charges for each atom or not, since those that do need to take the charge into account in the look-up table. While it may not lead to a significant difference in the evaluation results, we believe it is important to highlight this issue and we hope that the community will adopt a single definition for each metric going forward.
  2. 2.
    The stability metrics do not always provide a good reflection of chemical validity. As an example, the existing definition for atom stability allows an uncharged carbon atom with 3 (single bond equivalent) bonds to be considered ’stable’, but, to the best of our knowledge, this has little-to-no chemically valid basis. We believe this highlights the importance of using the metrics we have introduced for this task, energy and strain, since these will better reflect chemical soundness.
  3. 3.
    We have also found differences in the evaluation results for validity. When we re-evaluated EQGAT-diff on GEOM Drugs we found that our evaluation produced validity scores about 8% higher than those published in their paper. We have not been able to determine the cause of this but we have used the higher number in our comparisons.
We believe these findings underline the need for more in-depth evaluations of molecular generation models, as well as alignment within the community on the exact evaluation metrics to use. We hope that this discussion, as well as our introduction of energy and strain energy metrics, will begin to lay a foundation for more reproducible and coherent benchmarking.

### v3

@@ 6 RELATED WORK
In addition to the unconditional molecular generators we outlined above, a number of works have attempted to directly generate ligands within protein pockets (Peng et al., 2022; Guan et al., 2023; Schneuing et al., 2022; Cremer et al., 2024; Ziv et al., 2024). However, these models also suffer from the issues we outlined previously, including long-sampling times (100s - 1000s of seconds for 100 molecules (Schneuing et al., 2022)) and generating invalid chemical structures or molecules with very high strain energies (Harris et al., 2023; Buttenschoen et al., 2024). GraphBP (Liu et al., 2022), an autoregressive model for protein-conditioned generation, is able to generate ligands faster, but suffers from poor docking scores in comparison to more recent diffusion models.
@@
Outside of small molecule design flow-matching has recently gained traction with generative models for biomolecules. FoldFlow (Bose et al., 2024) and FrameFlow (Yim et al., 2023) are both recently introduced flow matching models for protein structure generation. Multiflow (Campbell et al., 2024) attempts to jointly generate protein sequence and structure and introduces the discrete flow models (DFM) framework for flow matching generation of discrete data. Stark et al. (2024b) also introduce a framework for flow matching on discrete data, DirichletFM, and apply this to DNA sequence design. Finally, Verma et al. (2023) use a conjoined system of ODEs to train a model to jointly generate antibody sequences and structures.
@@ 7 CONCLUSION
In this work we have presented Semla, a novel equivariant message passing architecture exhibiting significantly better efficiency and scalability than existing approaches 3D generation. We have further introduced SemlaFlow, a flow matching model for jointly generating the topology and 3D conformations of molecular graphs. SemlaFlow achieves state-of-the-art results on 3D molecular generation benchmarks with two orders-of-magnitude faster sampling times. We have also highlighted issues with current molecular generation evaluation metrics and proposed the use of energy per atom and strain energy per atom for evaluating the quality of generated molecular conformations.
@@
While we believe our model has made significant progress in solving key challenges for 3D molecular generators, many challenges remain. Firstly, the energies of the molecules generated by SemlaFlow are still significantly higher than that of the dataset; generating molecular coordinates with very high fidelity remains a problem for these models. Including further inductive biases or fine-tuning against an energy model could be an avenue to improve this in future work (Noé et al., 2019; Schreiner et al., 2023; Viguera Diez et al., 2024). Additionally, while SemlaFlow has shown significant efficiency improvements over existing methods, it still uses a fully-connected message passing component, limiting it’s scalability to larger molecular systems. We leave the further enhancement of the scalability of Semla to future work. We believe our model makes crucial step towards the practical application of 3D molecular generators, although we leave the integration of SemlaFlow into drug discovery workflows, either through RL-based fine-tuning or protein pocket conditioned generation, to future work.
Acknowledgements
@@
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Preliminary experiments were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS, project: 2024/22-33), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. The authors thank T. Le (Pfizer) for sharing code and weights of Le et al. (2024) ahead of publication.
Appendix A ABLATION EXPERIMENTS
This section includes additional experiments which were designed to explain the design decisions made in the Semla architecture and to assess the benefit of using Semla over other equivariant architectures for molecular generation.
A.1 Latent Message Size
Firstly, Table 4 assesses the impact of the $d_{l}$ hyperparameter which specifies the size of the latent dimension when performing latent attention. Each model was trained on GEOM Drugs for 100 epochs using the same training setup as the SemlaFlow model in Table 2.
While $d_{l}=256$ produces the best overall performance in terms of molecule stability, validity and energy, the model takes significantly longer to sample the same number of molecules. Crucially, the difference in performance between all four models is relatively small, despite the notable differences in sample time. For this reason we selected $d_{l}=64$ for all SemlaFlow models.
Table 4: Comparison of SemlaFlow models with different sizes of latent attention dimension $d_{l}$. All models are trained in identical conditions for 100 epochs. Sample time is measured by the average number of seconds to generate 1000 molecules. All metrics are averaged over 3 runs.
$d_{l}$ Mol Stab $\uparrow$ Valid $\uparrow$ Energy $\downarrow$ Strain $\downarrow$ Sample Time $\downarrow$
32 97.7±${}_{\pm\text{0.13}}$ 92.7±${}_{\pm\text{0.30}}$ 2.73±${}_{\pm\text{0.014}}$ 1.80±${}_{\pm\text{0.004}}$ 89.0
64 97.6±${}_{\pm\text{0.02}}$ 93.1±${}_{\pm\text{0.30}}$ 2.63±${}_{\pm\text{0.019}}$ 1.70±${}_{\pm\text{0.013}}$ 98.3
128 97.1±${}_{\pm\text{0.05}}$ 92.8±${}_{\pm\text{0.07}}$ 2.87±${}_{\pm\text{0.026}}$ 1.98±${}_{\pm\text{0.022}}$ 122.0
256 98.4±${}_{\pm\text{0.06}}$ 94.6±${}_{\pm\text{0.14}}$ 2.61±${}_{\pm\text{0.006}}$ 1.71±${}_{\pm\text{0.008}}$ 217.7
A.2 Architecture Ablation
We also wish to compare the generative performance and sampling efficiency of our proposed Semla architecture with existing E(3)-equivariant architectures for 3D molecular generation. To do this we train various network architectures within a consistent experimental setup – we apply the same flow matching training and inference procedure as SemlaFlow but swap out Semla for existing equivariant architectures. We benchmark Semla against the 4 layer and 9 layer versions of EGNN (Satorras et al., 2021) proposed in EDM (Hoogeboom et al., 2022), as well as the EQGAT network (Le et al., 2022) used in the state-of-the-art EQGAT-diff model Le et al. (2024). We also provide a comparison with an expanded, 16 layer version of EGNN, which has roughly the same number of parameters as Semla. In order to handle bond types, we modify the pairwise MLPs for EGNN on the first and last layer of the network in the same way as Semla. Each architecture is given a fixed training budget of 24 hours. Since existing models are not setup for self-conditioned inputs, all models, including Semla, are trained without self conditioning.
Table 5: Comparison of different architectures all trained on an identical flow matching setup. Sample time is measured by the average number of seconds to generate 1000 molecules. All metrics are averaged over 3 runs.
Architecture Mol Stab $\uparrow$ Valid $\uparrow$ Energy $\downarrow$ Strain $\downarrow$ Sample Time $\downarrow$
EGNN (4 layer) 71.4±${}_{\pm\text{0.76}}$ 55.1±${}_{\pm\text{0.06}}$ 10.89 ±${}_{\pm\text{0.059}}$ 10.03±${}_{\pm\text{0.063}}$ 69.3
EGNN (9 layer) 94.0±${}_{\pm\text{0.16}}$ 87.9±${}_{\pm\text{0.40}}$ 4.04±${}_{\pm\text{0.047}}$ 3.13±${}_{\pm\text{0.023}}$ 151.4
EGNN (16 layer) 94.7±${}_{\pm\text{0.06}}$ 89.9±${}_{\pm\text{0.06}}$ 3.97±${}_{\pm\text{0.079}}$ 3.07±${}_{\pm\text{0.065}}$ 532.0
EQGAT 97.1±${}_{\pm\text{0.07}}$ 83.9±${}_{\pm\text{0.33}}$ 4.19±${}_{\pm\text{0.013}}$ 3.32±${}_{\pm\text{0.019}}$ 337.4
Semla (Ours) 96.3±${}_{\pm\text{0.14}}$ 91.2±${}_{\pm\text{0.26}}$ 3.04±${}_{\pm\text{0.014}}$ 2.15±${}_{\pm\text{0.014}}$ 99.4
The results of this architecture ablation study are shown in Table 5. Semla shows better validities and significantly better energies than existing architectures, although has slightly lower molecule stability than EQGAT. Crucially though Semla is more than 3 times faster than EQGAT and more than 5 times faster than the similarly-sized 16 layer EGNN network while still producing molecules of comparable or higher quality.
Appendix B ADDITIONAL RESULTS ON GEOM DRUGS
In this section we include additional results of existing models on GEOM Drugs which we were not able to fit into the main text. We also include the results from Table 2 for full comparison. Many models have been proposed for this task recently and we chose to focus our comparison in the main text to models which learn to generate bonds rather than inferring them from coordinates.
Table 6 shows that while some models which infer bonds (shown in the top segment of the table) produce higher validities than those which do not, their atom and molecule stabilities are often significantly lower. However, since validity only measures whether a molecule can be sanitised by RDKit, we consider molecule stability to be a much better measure of the quality of the generated molecules. We include a further discussion and explanation for this in Appendix D.
Table 6: Additional results on GEOM Drugs, including models which infer bonds based on the generated coordinates. Since many of these models do not publish molecule stability results, numbers marked with * are estimates computed by taking $AS^{44}$ where $AS$ is atom stability and 44 is the average number of atoms in GEOM Drugs molecules. This follows the same procedure used in the EquiFM paper to estimate molecule stability.
Model Atom Stab $\uparrow$ Mol Stab $\uparrow$ Valid $\uparrow$ Unique $\uparrow$ Novel $\uparrow$ NFE
EDM 81.3 0.0* 1000
GCDM 89.0 5.2 1000
MUDiff 84.0 60.9 98.9 1000
GFMDiff 86.5 3.9 500
EquiFM 84.1 0.0* 98.9
GeoBFN 86.2 0.0* 91.7 2000
GeoLDM 98.9 61.5* 99.3 1000
FlowMol 99.0 67.5 51.2 100
MiDi 99.8 91.6 77.8 100.0 100.0 500
EQGAT-diff 99.8±${}_{\pm\text{0.0}}$ 93.4±${}_{\pm\text{0.21}}$ 94.6±${}_{\pm\text{0.24}}$ 100.0±${}_{\pm\text{0.0}}$ 99.9±${}_{\pm\text{0.07}}$ 500
SemlaFlow (Ours) 99.8±${}_{\pm\text{0.0}}$ 97.3±${}_{\pm\text{0.08}}$ 93.9±${}_{\pm\text{0.19}}$ 100.0±${}_{\pm\text{0.0}}$ 99.6±${}_{\pm\text{0.03}}$ 100
Appendix C FURTHER MODEL AND TRAINING DETAILS
This section provides further detail on the design of the full Semla architecture, as well further training details including the hyperparameters used to trian QM9 and GEOM Drugs models.
C.1 The Semla Architecture
As mentioned in Section 3, we do not carry pairwise edge features through the model, but rather encode the edge information into the node features on the first layer and then generate edge features on the final layer. After these initial edge features are generated they are passed through a learnable refinement component where the final predicted coordinates are also given as input, along with the final invariant node features. The remaining components of the model are used to encode atom and bond types and predict distributions for atom types, bond types and formal charges. A full overview of a Semla model is shown in Figure 3.
Figure 3: An overview of a full Semla model. A Semla model is created from a stack of Semla layers where the first layer encodes bond information into the node features and the final layer generates bond information from the pairwise features.
C.2 Training Details
@@
All models were trained with the AMSGrad (Reddi et al., 2018) variant of the Adam optimiser (Kingma and Ba, 2014) with a learning rate (LR) of 0.0003. We also apply linear LR warm-up, using 2000 warm-up steps for QM9 and 10000 warm-up steps for GEOM Drugs. During training we clip the norms of the gradients at 1.0 for all models. Loss weightings $(\lambda_{x},\lambda_{a},\lambda_{b},\lambda_{c})=(1.0,0.2,0.5,1.0)$ were used to train QM9 models. The same weightings were used for GEOM Drugs, except we set $\lambda_{b}=1.0$.
@@
In order to make the training as efficient as possible we place molecules in the dataset into buckets based on their size, and then form minibatches for training and evaluation within the buckets. This ensures that all batches contain similarly sized molecules so that the amount of padding within each batch is minimised. With this setup we can also apply a cost function to select the batch size for each bucket separately; since the amount of memory required to process a molecule increases quadratically with the number of atoms, this helps to balance the GPU memory consumption for each batch. In practice, though, we simply apply a linear cost function and use a batch size of 4096 atoms per batch for all SemlaFlow models. While we have found our bucketing scheme leads to a significant increase in training speed, it may also introduce additional bias into the training since molecules within each batch are no longer selected completely at random. Although we have not attempted to quantify this bias our results seem to show that bucketing is not significantly detrimental to performance.
Appendix D EVALUATION METRICS
@@
  • Validity is given by the proportion of molecules which can be successfully sanitised using RDKit.
  • Uniqueness measures the proportion of generated molecules which are unique. This is calculated by comparing molecules based on their canonical SMILES representation.
  • Novelty measures the proportion of generated molecules which are not in the training set.
D.1 Issues with Existing Metrics
While a lot of progress has been made recently in improving benchmarks for structure-based drug design (Harris et al., 2023; Buttenschoen et al., 2024), we believe many evaluation metrics used for unconditional 3D generation are still not fit for purpose. We hope that the introduction of energy per atom and strain energy per atom will help to alleviate this but believe it is still important to discuss the shortcomings of existing evaluation methods. In this section we highlight issues with the standard benchmark metrics for this task.
Validity
Since RDKit is free to add implicit hydrogens and, in some cases, to modify the formal charge on atoms, RDKit validity tells us very little about the "correctness" of a molecule. This is shown explicitly in Figure 4; all molecules shown in the figure would be marked as valid by RDKit, even though only the rightmost molecule has the correct number of bonds. Additionally, the validity metric on its own is not capable of flagging disconnected molecules – generated molecules which contain multiple fragments which are not connected by any bond. The validity metric is, however, capable of spotting when a molecule has too many bonds.
Atom and Molecule Stability
Different models sometimes use different lookup tables to compute atom (and therefore molecule) stability, potentially leading to different evaluation results for the same generated molecules. Often this depends on whether models predict formal charges for each atom or not, since those that do need to take the charge into account in the look-up table. Additionally, we have found that existing lookup tables used to define atom stability have little basis in chemical validity. As an example, the existing definition for atom stability allows an uncharged carbon atom with 3 (single covalent bond equivalent) bonds to be considered ’stable’, although this has little-to-no chemically valid basis.
Conclusion
We believe the issues highlighted here reflect the importance of using metrics such as energy which is able to quantify the quality of both the generated topology as well as the generated molecular conformation. We also hope that highlighting these issues will lead to more in-depth evaluations of 3D unconditional molecular generation models in the future and lay a foundation for more reproducible and coherent benchmarking.
Figure 4: Examples of molecules which would be considered valid by RDKit.